Fill in the blanks with real numbers.
(11 - 7i)(2 + i) = ___ + ___ *i

Answer by math_tutor2020(3816)


(11-7i)*(2+i)
= x*(2+i)
= x*2+x*i
= (11-7i)*2+(11-7i)*i
= 22-14i+11i-7i^2
= 22-14i+11i-7(-1)
= 22-14i+11i+7
= (22+7)+(-14i+11i)
= 29-3i

Therefore,
(11-7i)*(2+i) = 29-3i

Verification using WolframAlpha